# Learning large softmax mixtures with warm start EM

### v1

@@ 2.3 MoM estimation of the softmax mixture parameters
Grant Assumptions 2, 3 and 4. There exists a positive constant $D$ depending on $K$, $B$, $\Delta_{1}$, and $\underline{\alpha}$ such that, up to relabeling, it holds
$\|\bm{\alpha}^{*}-\widehat{\bm{\alpha}}\|_{2}$ $\leq D\|\bm{m}-\widetilde{\bm{m}}\|_{2},$
$\|{\bm{\theta}}_{k}-\widehat{{\bm{\theta}}}_{k}\|_{2}$ $\leq D\left(L\|\bm{m}-\widetilde{\bm{m}}\|^{2}_{2}+\sum_{i=2}^{L}\|\bm{m}_{1;i}-\widehat{\bm{m}}_{1;i}\|^{2}_{2}\right)^{1/2},\quad\quad\text{for all $k\in\{1,\dots,K\}$.}$
The following proposition gives the error rates of the softmax mixture parameters. It is an immediate consequence of Propositions 2 and 3, and of Corollary 1.
Proposition 4.
Adopt the same assumptions as Proposition 2 and Proposition 3. Then, a.s. $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$, we have
$\lim_{C\to\infty}\limsup_{N\to\infty}\mathbb{P}\left\{\|\bm{\alpha}^{*}-\widehat{\bm{\alpha}}\|_{2}\geq C\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right\}$ $=0,$
$\lim_{C\to\infty}\limsup_{N\to\infty}\mathbb{P}\left\{\|{\bm{\theta}}_{k}-\widehat{{\bm{\theta}}}_{k}\|_{2}\geq C\sqrt{L}\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right\}$ $=0,$
for each $k\in[K]$, and any $p\geq 1$.
Remark 4.
Although we have stated Proposition 3 without explicit constants, the dependence on $K$, $B$, $\Delta_{1}$, and $\underline{\alpha}$ can be extracted from the proof. In particular, see for instance display (65), the dependence on $\Delta_{1}$ is poor: the constant $D$ can be shown to scale as $\Delta_{1}^{-cK}$ for some absolute constant $c$. While it is possible that the exponent can be improved, the exponential dependence of this constant on $K$ cannot be entirely avoided, even for univariate mixtures ($L=1$). This follows from the fact that, when $L=1$, for any $K\geq 2$ and sufficiently small $\Delta>0$, there exist a pair of $K$-atomic probability measures $\rho$ and $\rho^{\prime}$ on $[-1,1]$ with support $\{{\bm{\theta}}_{1},\dots,{\bm{\theta}}_{K}\}$ and $\{{\bm{\theta}}^{\prime}_{1},\dots,{\bm{\theta}}^{\prime}_{K}\}$, all of which are separated by at least $\Delta$, and such that
$\|{\bm{\theta}}^{\prime}_{k}-{\bm{\theta}}_{k}\|_{2}\geq C_{K}\Delta^{-2K+2}\|\bm{m}-\bm{m}^{\prime}\|_{2},\quad\quad\text{for all $k\in\{1,\dots,K\}$}$ (28)
where $\bm{m}$ and $\bm{m}^{\prime}$ are the vectors of the first $12K-$ 1 moments of $\rho$ and $\rho^{\prime}$, respectively. We justify this fact in Section A.6. This example shows that any deterministic bound on the distance between the atoms in terms of the moment difference for $K$-atomic distributions with well separated atoms must involve a prefactor of the same type as appears in (28). Since Proposition 3 is a bound of this type, we conclude that the $\Delta_{1}^{-cK}$ scaling of $D$ is essentially tight.
Remark 5.
The parametric-type rates of Proposition 4 hold when the mixture atoms are well separated. In particular, the proof of Proposition 3 reveals that its conclusion holds when each atom of $\widetilde{\rho}_{1}$ and $\rho$ is at least $\Delta_{1}/2$ away from all but $\ell^{\prime}=1$ other atom (itself). However, if $\ell^{\prime}>1$, Proposition 4 in Wu and Yang (2020) shows that we cannot expect a parametric rate in the estimation of ${\bm{\theta}}_{k}$, even in one dimension, as display (64) in the proof then becomes
$W_{1}(\widetilde{\rho}_{1},\rho_{1})\leq 2K\left({2K4^{2K-1}2^{2K-\ell^{\prime}-1}\over\Delta_{1}^{2K-\ell^{\prime}-1}}\right)^{1\over\ell^{\prime}}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}^{1\over\ell^{\prime}},$ (29)
a rate that will be inherited by $|\theta_{1k}-\widehat{\theta}_{1k}|$, for each $k$, via (61). In the worst case, when $\ell^{\prime}=2K-1$, we obtain $W_{1}(\widetilde{\rho}_{1},\rho_{1})\lesssim K\|\widetilde{\bm{m}}-\bm{m}\|_{2}^{1/(2K-1)},$ by Proposition 1 in Wu and Yang (2020). Thus, although consistent estimation of the softmax mixture parameters will continue to hold when the atoms are distinct, but not well separated, neither the estimation of ${\bm{\theta}}_{k}$ nor that of ${\bm{\alpha}}$ can be expected to follow a parametric decay rate. This is confirmed by our simulation results in Section 4.
Nevertheless, despite their potential shortcomings, the MoM estimates remain in ball of appropriate radius from the true parameters, making them excellent starting values for the EM algorithm described and analyzed in Section 3, where we make this precise.
2.4 Extension to $\mu=\mathcal{N}_{L}(0,\Sigma)$
When $X\sim\mathcal{N}_{L}(0,\Sigma)$, one can continue to use the estimates (14) and (18) relative to the functions $h_{r}$ and $h_{r1;i}$ given by (19), noting that they estimate moments and mixed-moments of the scaled random variable $\Sigma Z$, for the latent variable $Z\sim\rho$ introduced in Section 2.1. To see this, it is enough to consider $r$-th moments, the argument for mixed moments is identical.
Since $X\sim\mathcal{N}_{L}(0,\Sigma)$, then $U:=\Sigma^{-1/2}X\sim\mathcal{N}_{L}(0,\mathbf{I}_{L})$. Consider the left hand side of (21), with $h_{r}$ given by (19). Then, for generic ${\bm{\theta}}\in\mathbb{R}^{L}$, and given $v\in\mathbb{R}^{L}$
$\mathbb{E}\left[H_{r}(X^{\top}v)\exp(X^{\top}{{\bm{\theta}}})\right]\over\mathbb{E}\left[\exp(X^{\top}{{\bm{\theta}}})\right]$ $=$ ${\mathbb{E}\left[H_{r}(U^{\top}u)\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]\over\mathbb{E}\left[\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]}\ =\ (u^{\top}\bar{{{\bm{\theta}}}})^{r}\ =\ (v^{\top}\Sigma{{\bm{\theta}}})^{r}$ (30)
with $u:=\Sigma^{1/2}v$ and $\bar{{{\bm{\theta}}}}:=\Sigma^{1/2}{{\bm{\theta}}}$, where the second equality holds by Proposition 1, since $U$ is a standard Gaussian on $\mathbb{R}^{L}$. This shows that the left hand side of $\bar{m}_{r}$ in (16) has the a.s. $\mu=\mathcal{N}_{L}(0,\Sigma)$ limit equal to the $r$-th moment of the first coordinate of the random vector $\bar{Z}\in\mathbb{R}^{L}$ where $\bar{Z}$ takes values in $\{R^{\top}\Sigma{\bm{\theta}}_{1}^{*},\ldots,R^{\top}\Sigma{\bm{\theta}}_{K}^{*}\}$ with probability masses in ${\bm{\alpha}}^{*}$. Therefore, all the results established in the preceding subsections, continue to hold, with respect to $\mu=\mathcal{N}_{L}(0,\Sigma)$, with the difference that now $\widehat{{\bm{\theta}}}_{k}$ constructed above estimates $R^{\top}\Sigma{{\bm{\theta}}}_{k}^{*}$, for each $k\in[K]$, and thus Proposition 4 above applies to $\|\widehat{\bm{\theta}}_{k}-R^{\top}\Sigma{{\bm{\theta}}}_{k}^{*}\|_{2}$, after we replace Assumption 4 by
Assumption 5.
There exits $\Delta_{2}>0$ such that
$\Delta(v^{\top}\Sigma{\bm{\theta}}_{1}^{*},\ldots,v^{\top}\Sigma{\bm{\theta}}_{K}^{*})\geq\Delta_{2}.$
An estimator of ${{\bm{\theta}}}_{k}^{*}$, for all $k\in[K]$, can be immediately constructed as
$\widehat{\bm{\theta}}_{k}^{*}:=\widehat{\Sigma}^{-1}R\widehat{{\bm{\theta}}}_{k},$
for an estimator $\widehat{\Sigma}$ of $\Sigma$. Since we work in the regime $L<p$, it is enough to consider the sample covariance matrix
$\widehat{\Sigma}={1\over p}\mathbf{X}^{\top}\mathbf{X}={1\over p}\sum_{i=1}^{p}X_{i}X_{i}^{\top}.$ (31)
We make the following assumption on the spectrum of the covariance matrix $\Sigma$.
Assumption 6.
Under 6 and $L\leq p$, existing results guarantee that
$\mathbb{P}\left\{\|\widehat{\Sigma}-\Sigma\|_{\mathrm{op}}~{}\gtrsim~{}\overline{\sigma}^{2}\sqrt{L+\log(p)\over p}\right\}\leq 2p^{-2}.$ (32)
See, for instance, Vershynin (2018). Then, the estimation rate of $\widehat{{\bm{\theta}}}_{k}^{*}$ will inherit that given by Proposition 4, plus a term that reflects the estimation of $\Sigma$. We sketch the standard argument below, informally. For every $k\in[K]$ we have, with high probability,
$\|\widehat{{\bm{\theta}}}_{k}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}$$\leq$ $\|\widehat{\Sigma}^{-1}-\Sigma^{-1}\|_{\mathrm{op}}\|\widehat{\bm{\theta}}_{k}\|_{2}+\|\Sigma^{-1}\|_{\mathrm{op}}\|\widehat{\bm{\theta}}_{k}-R^{\top}\Sigma{\bm{\theta}}_{k}^{*}\|_{2}$
$\lesssim$ $\|\widehat{\Sigma}-\Sigma\|_{\mathrm{op}}\|\Sigma{\bm{\theta}}_{k}^{*}\|_{2}+\|\widehat{\bm{\theta}}_{k}-R^{\top}\Sigma{\bm{\theta}}_{k}^{*}\|_{2}.$
We summarize the concluding result below.
Corollary 2.
Grant Assumptions 2, 3, 5 and 6. Then, a.s $\mu=\mathcal{N}(0,\Sigma)$, we have
$\lim_{C\to\infty}\limsup_{N\to\infty}\mathbb{P}\left\{\|\widehat{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}\geq C\sqrt{L}\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}+\sqrt{\frac{\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right\}$ $=0.$
2.5 Practical choice of the primary axis
As mentioned in Remark 4, finding a good primary axis $v$ relative to which $\Delta_{1}$ of Assumption 4 is large is crucial for the success of the MoM estimation technique. For $\mu=\mathcal{N}_{L}(0,\Sigma)$, this gets translated to $\Delta_{2}$ in 5. In this section, we treat directly the general case $\mu=\mathcal{N}_{L}(0,\Sigma)$ and offer a practical recommendation for choosing a unit vector $v\in\mathbb{S}^{L-1}$ such that the projections ${v}^{\top}\Sigma{\bm{\theta}}_{1}^{*},\dots,{v}^{\top}\Sigma{\bm{\theta}}_{K}^{*}$ are well separated.
In rare cases, the statistician may have a priori knowledge of a good direction $v$ for which a lower bound $\Delta_{2}$ on $\Delta(v^{\top}\Sigma{\bm{\theta}}_{1}^{*},\dots,v^{\top}\Sigma{\bm{\theta}}_{K}^{*})$, defined in (7), is sufficiently large. In general, to obtain results that hold uniformly over the parameter space, one could choose $v$ randomly on the sphere $\mathbb{S}^{L-1}$. Indeed, consider the following separation condition on ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$.
Assumption 7.
There exists a sufficiently large constant $C_{\Delta}$ such that
$\Delta^{2}:=\min_{k\neq k^{\prime}}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2}\geq C_{\Delta}$ (33)
Then, with $\underline{\sigma}^{2}$ defined in 6, a simple probabilistic argument (see, Lemma 19) gives that, for any $t\in(0,1)$ and any $v$ uniformly drawn from $\mathbb{S}^{L-1}$,
$\mathbb{P}\left\{\Delta(v^{\top}\Sigma{\bm{\theta}}_{1}^{*},\dots,v^{\top}\Sigma{\bm{\theta}}_{K}^{*})\geq{\underline{\sigma}^{2}C_{\Delta}\over K^{2}\sqrt{L}}~{}t\right\}\geq 1-t.$ (34)
Recall that in our setting we view $K$ as being a relatively small integer. The above display gives the following order of magnitude of $\Delta_{2}$ as
$\Delta_{2}\asymp{1\over K^{2}\sqrt{L}}.$
Following an identical argument to that of the proof of Proposition 3, applied to the re-scaled targets $R^{\top}\Sigma{{\bm{\theta}}}_{k}^{*}$, for each $k\in[K]$, the constant $D$ scales as multiples of $\Delta_{2}^{-cK}$, for some $c>0$. Thus, in an application in which $L<K$, it will scale as $K^{c^{\prime}K}$, for some $c^{\prime}>0$. In most of the contemporary applications mentioned in the Introduction, however, $K<L$, and typically $K\ll L$, as in most natural language applications. Then, the constant $D$ will scale as $L^{c^{\prime\prime}K}$, for some $c{{}^{\prime\prime}}>0$. Fortunately, it is possible to eliminate the dependency of $D$ on $L$ altogether, by choosing a direction $v$ from a lower dimensional space, which will allow us to improve upon (34). This will be done via the dimension reduction step below. A similar dimension reduction technique is also used in density estimation for multivariate Gaussian mixture models (Doss et al., 2023).
The basic idea is to first estimate the span of $\Sigma{\bm{\theta}}_{1}^{*},\ldots,\Sigma{\bm{\theta}}_{K}^{*}$ which is at most a $K$-dimensional subspace of $\mathbb{R}^{L}$, and then sample the unit vector $v$ uniformly in this subspace. Specifically, we propose to use the first $K$ eigenvectors of
@@ 2.5 Practical choice of the primary axis
$\widehat{\Gamma}:={1\over N}{\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{\Sigma}={1\over N}\sum_{\ell=1}^{N}Y_{\ell}Y_{\ell}^{\top}-{1\over p}\sum_{j=1}^{p}X_{j}X_{j}^{\top}$ (35)
to estimate the span of $\Sigma{\bm{\theta}}_{1}^{*},\ldots,\Sigma{\bm{\theta}}_{K}^{*}$. The following proposition provides its justification and its proof, stated in Section A.7, reasons similarly as in Propositions 1 and 2. While we generally consider the ambient dimension $L$ to be fixed, it is allowed to grow with the sample sizes, under the following growth condition.
Assumption 8.
There exist constants $\delta>0$ and $c>0$ such that
$L$ $\leq c~{}\min\left({p^{1-\delta}}~{},~{}\sqrt{N\over\log N}~{},~{}{N\over\log N\log p}\right).$ (36)
Proposition 5.
Grant Assumptions 2, 6 and 8. Let $\Gamma=\sum_{k=1}^{K}\alpha_{k}^{*}\Sigma{\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}\Sigma.$ Then one has
@@
$\mathbb{P}\left\{\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{(L+\log(p))\log(N)\over N}+\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}~{}\Big{|}~{}X_{1},\ldots,X_{p}\right\}\geq 1-N^{-1}$
$\mu$-almost surely, as $p\to\infty$.
Collect the first $K$ eigenvectors of $\widehat{\Gamma}$ in the matrix $\widehat{V}\in\mathbb{O}_{L\times K}$. In view of (5), we propose to choose the projection vector $v$ as
@@
where the vector $u\in\mathbb{R}^{L}$ contains i.i.d. entries of $\mathcal{N}(0,1)$. The following proposition gives a lower bound on the desired minimum pairwise separation relative to this choice for $v$. Its proof is given in Section A.8. It is worth mentioning that our analysis does not require any spectral condition on $\Gamma$.
Proposition 6.
Grant Assumptions 2, 3, 6, 7 and 8. Then for any $t\in(0,1)$ and any $v$ drawn as (37), one has
$\mathbb{P}\left\{\Delta(v^{\top}\Sigma{\bm{\theta}}_{1}^{*},\ldots,v^{\top}\Sigma{\bm{\theta}}_{K}^{*})\geq{\underline{\sigma}^{2}C_{\Delta}\over 2K^{2}\sqrt{K}}~{}t\right\}\geq 1-t.$
Comparing to (34), using the dimension reduction in (37) will allow us to take
$\Delta_{2}\asymp{1\over K^{2}\sqrt{K}},$
thereby eliminating the dependency on $L$ in the constant $D$ of Proposition 3, provided that Assumptions 7 and 8 hold.
@@
In practice we recommend to take multiple random projection vectors $\{v_{1},\ldots,v_{n}\}$, and select the one that yields the largest separation. However since the separation $\Delta(v_{i}^{\top}\Sigma{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}\Sigma{\bm{\theta}}_{K}^{*})$ is unknown, we propose to use the following criterion. For any $i\in[n]$, compute the moment vector $\widehat{\bm{m}}(v_{i})$ as in (14) with $h_{r}$ given by (19), for each $v_{i}$, and its denoised version $\widetilde{\bm{m}}(v_{i})$ as in (24), form the moment matrix of $\widetilde{\bm{m}}(v_{i})$ as
$\widetilde{\bm{M}}(v_{i}):=\begin{pmatrix}1&\widetilde{m}_{1}(v_{i})&\dots&\widetilde{m}_{K-1}(v_{i})\\ \widetilde{m}_{1}(v_{i})&\widetilde{m}_{2}(v_{i})&\dots&\widetilde{m}_{K}(v_{i})\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}(v_{i})&\widetilde{m}_{K}(v_{i})&\dots&\widetilde{m}_{2K-2}(v_{i})\end{pmatrix},$
@@
$i_{*}=\operatorname*{argmax}_{i\in[n]}~{}\det\left(\widetilde{\bm{M}}(v_{i})\right).$
@@
$\det({\bm{M}}(v_{i}))={1\over K!}\prod_{1\leq k<k^{\prime}\leq K}\left(v_{i}^{\top}\Sigma{\bm{\theta}}_{k}^{*}-v_{i}^{\top}\Sigma{\bm{\theta}}_{k^{\prime}}^{*}\right)^{2}$

### v2

@@ 3.4 Subspace estimation via MoM under softmax mixtures
Grant Assumptions 2 and 3. Assume
$\|(\mu(x))^{-1}\nabla^{2}\mu(x)\|_{\mathrm{op}}\leq C\|x\|_{2}^{2},\qquad\forall~x\in\mathbb{R}^{L}$ (55)
for some constant $C>0$. Then for any converging sequence $\epsilon_{p}^{\prime}$, on the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime}):=\left\{\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}\leq\epsilon_{p}^{\prime}\right\},$ for sufficiently large $p$, with probability at least $1-N^{-1}$, one has
@@
$\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}~\leq~C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N\over N}+\epsilon_{p}^{\prime}.$
As mentioned in Remark 7, condition (55) is a mild Lipschitz requirement, which holds, for instance, for $\mu$ being Gaussian, or for finite Gaussian mixtures.
In the setting of Theorem 3 for fixed $\mu$, it is easy to see that $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$ holds $\mu$ almost surely, with $\epsilon_{p}^{\prime}\leq\epsilon_{p}\sqrt{L}=\mathcal{O}(\sqrt{\log\log p/p})$ for sufficiently large $p$.
For $\mu$ allowed to depend on $p$, we give explicit results for Gaussian below. Similar results hold for sub-Gaussian $\mu$ under conditions mentioned in Remark 7.
Example 3.
For $\mu=\mathcal{N}_{L}(0,\Sigma)$, we prove in Section C.10 that the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$ holds for $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{L\log(p)/p})$ with probability at least $1-p^{-1}$, provided that $p\geq L^{a}$ for some $a>3$. On the other hand, using (2), the choice of $\widehat{\Gamma}$ in (54) becomes
$\widehat{\Gamma}={1\over N}\sum_{\ell=1}^{N}\Sigma^{-1}Y_{\ell}Y_{\ell}^{\top}\Sigma^{-1}-\Sigma^{-1}.$ (56)
As a result, with probability at least $1-N^{-1}-p^{-1}$, one has
@@
$\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}~\lesssim~\lambda_{L}^{-1}(\Sigma)\sqrt{(L+\log(p))\log(N)\over N}+\lambda_{L}^{-1}(\Sigma)\sqrt{L\log(p)\over p}.$ (57)
When $\Sigma$ is unknown, it can be consistently estimated by the sample $\widehat{\Sigma}=p^{-1}\mathbf{X}^{\top}\mathbf{X}$.
3.4.1 Application to practical choice of the primary axis
Recall in Section 3.1 that we have chosen the primary axis as $\mathbf{e}_{1}$, which leads to $\Delta_{1}$ in 4. As mentioned in Remark 10, finding a good primary axis $v$ relative to which $\Delta_{1}$ is large is crucial for the success of the MoM estimation technique. In rare cases, the statistician may have a priori knowledge of a good direction $v$ for which a lower bound on $\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\dots,v^{\top}{\bm{\theta}}_{K}^{*})$, defined in (28), is sufficiently large. In general, to obtain results that hold uniformly over the parameter space, one could choose $v$ randomly on the sphere $\mathbb{S}^{L-1}$. Let
$\Delta^{2}:=\min_{k\neq k^{\prime}}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2},$ (58)
then a simple probabilistic argument (see, Lemma 32) gives that, for any $t\in(0,1)$ and any $v$ uniformly drawn from $\mathbb{S}^{L-1}$,
$\mathbb{P}\left\{\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\dots,v^{\top}{\bm{\theta}}_{K}^{*})\geq{t\Delta\over K^{2}\sqrt{L}}\right\}\geq 1-t.$ (59)
Fix such $v$, and let $R=(v,w_{2},\ldots,w_{L})$ be an $L\times L$ rotation matrix with orthonormal columns. Following an identical argument to that of the proof of Proposition 2, applied to the re-scaled targets $R^{\top}{\bm{\theta}}_{1}^{*},\ldots,R^{\top}{\bm{\theta}}_{K}^{*}$, the constant $D$ scales as multiples of $(\Delta/(K^{2}\sqrt{L}))^{-cK}$, for some $c>0$. Thus, it scales as $K^{\mathcal{O}(K)}$ if $L<K$, while it scales as $L^{\mathcal{O}(K)}$ otherwise. Fortunately, it is possible to eliminate the dependency of $D$ on $L$ altogether, by choosing a direction $v$ from the lower-dimensional subspace spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, which will allow us to improve upon (59). Recall that the subspace of ${\bm{\theta}}_{k}^{*}$’s are contained in $\Gamma$ given by (52). Denote by $\widehat{V}\in\mathbb{O}_{L\times K}$ the first $K$ eigenvectors of its estimator $\widehat{\Gamma}$ in (54). In view of Proposition 3, we propose to choose the projection vector $v$ as
@@ 3.4.1 Application to practical choice of the primary axis
where the vector $u\in\mathbb{R}^{L}$ contains i.i.d. entries of $\mathcal{N}(0,1)$. The following lemma gives a lower bound on the desired minimum pairwise separation relative to this choice for $v$. It is worth mentioning that our analysis does not require any spectral condition on $\Gamma$.
Lemma 2.
Grant Assumptions 2, 3, 5 and condition (55). Then for any $t\in(0,1)$ and any $v$ drawn as (60), on the event $C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N/N}+\epsilon_{p}^{\prime}\leq\underline{\alpha}\Delta^{2},$ one has
$\mathbb{P}\left\{\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\ldots,v^{\top}{\bm{\theta}}_{K}^{*})\geq{t\Delta\over 2K^{2}\sqrt{K}}\right\}\geq 1-t.$
Proof.
The proof is given in Section C.11. ∎
Compared to (59), the dimension reduction in (60) eliminates the dependency on $L$ in the constant $D$ of Proposition 2.
@@
In practice we recommend to take multiple random projection vectors $\{v_{1},\ldots,v_{n}\}$, and select the one that yields the largest separation. However since the separation $\Delta(v_{i}^{\top}{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}{\bm{\theta}}_{K}^{*})$ is unknown, we propose to use the following criterion. For any $i\in[n]$, compute the moment vector $\widehat{\bm{m}}(v_{i})$ as in (49) with $h_{r}$ given by (124), for each $v_{i}$, and its denoised version $\widetilde{\bm{m}}(v_{i})$ as in (43), form the moment matrix of $\widetilde{\bm{m}}(v_{i})$ as
$\widetilde{\mathbf{M}}(v_{i}):=\begin{pmatrix}1&\widetilde{m}_{1}(v_{i})&\dots&\widetilde{m}_{K-1}(v_{i})\\ \widetilde{m}_{1}(v_{i})&\widetilde{m}_{2}(v_{i})&\dots&\widetilde{m}_{K}(v_{i})\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}(v_{i})&\widetilde{m}_{K}(v_{i})&\dots&\widetilde{m}_{2K-2}(v_{i})\end{pmatrix},$
@@
$i_{*}=\operatorname*{argmax}_{i\in[n]}~\det\left(\widetilde{\mathbf{M}}(v_{i})\right).$
@@
$\det({\mathbf{M}}(v_{i}))={1\over K!}\prod_{1\leq k<k^{\prime}\leq K}\left(v_{i}^{\top}{\bm{\theta}}_{k}^{*}-v_{i}^{\top}{\bm{\theta}}_{k^{\prime}}^{*}\right)^{2}$